Blood is flowing through an artery of radius 2 mm at a rate of 40 cm/s. Determine the flow rate and the volume that passes through the artery in a period of 30 s.


Round your answer to the nearest 0 decimal place.


Answers:

  • flow rate: cm3/s
  • volume: cm3
1
Expert's answer
2021-04-29T10:45:35-0400

Cross-section area of artery, "A=\pi r^2=3.14\times(2\times10^{-1})^2"

"A=0.1256\space cm^2"

Velocity of flow, "v=40\space cm/s"

Volumetric Flow rate, "q=Av=0.1256\times40=5.024\space cm^3/s"

Volume passing through artery in 30s "=qt=5.024\times30=150.72\space cm^3"